# Integrating Multi-scale Contextualized Information for Byte-based Neural Machine Translation

### v1

@@ Abstract
Subword tokenization is a common method for vocabulary building in Neural Machine Translation (NMT) models. However, increasingly complex tasks have revealed its disadvantages. First, a vocabulary cannot be modified once it is learned, making it hard to adapt to new words. Second, in multilingual translation, the imbalance in data volumes across different languages spreads to the vocabulary, exacerbating translations involving low-resource languages. While byte-based tokenization addresses these issues, byte-based models struggle with the low information density inherent in UTF-8 byte sequences. Previous works enhance token semantics through local contextualization but fail to select an appropriate contextualizing scope based on the input. Consequently, we propose the Multi-Scale Contextualization (MSC) method, which learns contextualized information of varying scales across different hidden state dimensions. It then leverages the attention module to dynamically integrate the multi-scale contextualized information. Experiments show that MSC significantly outperforms subword-based and other byte-based methods in both multilingual and out-of-domain scenarios. Code can be found in https://github.com/ictnlp/Multiscale-Contextualization.
@@ 1 Introduction
However, a feature of UTF-8 encoding hinders conventional Transformer model Vaswani et al. (2017) from adapting well to byte-based vocabulary: a single character may correspond to 1 to 4 UTF-8 bytes. The number is 1 for English characters, but Arabic and many Asian languages require multiple bytes to represent a single character. Therefore, sometimes a single byte does not have a determined meaning; it requires the integration of local information to encode its semantics. To address that, various methods have been proposed for integrating local contextual information. MEGABYTE segments a sentence into blocks of 4 and simply concatenates the tokens Yu et al. (2023). CHARFORMER segments a sentence 4 times with block sizes ranging from 1 to 4 each, and employs mean-pooling to perform local integration. The weighted-sum of 4 results yields the final output Tay et al. (2022). LOBEF proposed Byte-nCF, replacing mean-pooling with Convolutional Neural Networks (CNNs) for better performance Sreedhar et al. (2023).
@@ 3.1 Datasets
We use a multilingual TED corpus of 59 languages Qi et al. (2018), TED-59, which includes both high and low-resource languages. All cases are English-centered. We collect the raw data from Salesky et al. (2023) and preprocess it with two subword-level vocabularies and a byte-level one. For subword-based baseline system, we leverage SentencePiece Kudo and Richardson (2018) and train a 32k vocabulary on the training set. We also incorporate the 250k mBART-50 Liu et al. (2020); Tang et al. (2020) vocabulary for full lexical coverage. For Byte-level systems, we preprocess data with a 256 vocabulary using scripts111https://github.com/UriSha/EmbeddinglessNMT/blob/master/embeddingless_scripts/byte_preprocess.sh from Shaham and Levy (2021).
@@ 3.2 Models
Byte-nCF Sreedhar et al. (2023): A strong byte-based model performing well under low-resource settings. The structure hyper-parameters are of default setting222https://github.com/makeshn/LOBEF_Byte_NMT/blob/main/embeddingless_scripts/train_byte_ncf.sh.
@@ 4.4 Contextualization Scales
For the selection of series, we experiment the smaller scales "0,0,1,1,3,3,5,5", the larger scales "0,0,1,1,5,5,7,7", and the balanced scales "0,0,1,1,3,5,5,7".
The results are exhibited in Table 4. First, the model of balanced scales performs the best averagely, because it is provided with contextualized information of more scales and has more options to choose. Second, if we dismiss the balanced one and compare the other models, the performance is related to the language groups. The sign indicates the better performance between the models. For "Byte-1" group, a small scale is sufficient to model certain semantics, so the smaller scaled model performs better. For "Byte-2" and "Byte-3" groups, it requires larger contextual scales to form certain meanings. As a result, the larger scaled model performs better.
These discoveries shade light on the selection of series, which needs to be compatible with the language of input text.
| series | 00113355 | 00113557 | 00115577 |
| Byte-1 | 30.190* | 30.219 | 29.941 |
| Byte-2 | 21.251 | 21.545 | 21.601* |
| Byte-3 | 10.302 | 10.712 | 10.686* |

### v2

@@ Abstract
Subword tokenization is a common method for vocabulary building in Neural Machine Translation (NMT) models. However, increasingly complex tasks have revealed its disadvantages. First, a vocabulary cannot be modified once it is learned, making it hard to adapt to new words. Second, in multilingual translation, the imbalance in data volumes across different languages spreads to the vocabulary, exacerbating translations involving low-resource languages. While byte-based tokenization addresses these issues, byte-based models struggle with the low information density inherent in UTF-8 byte sequences. Previous works enhance token semantics through local contextualization but fail to select an appropriate contextualizing scope based on the input. Consequently, we propose the Multi-Scale Contextualization (MSC) method, which learns contextualized information of varying scales across different hidden state dimensions. It then leverages the attention module to dynamically integrate the multi-scale contextualized information. Experiments show that MSC significantly outperforms subword-based and other byte-based methods in both multilingual and out-of-domain scenarios. We have uploaded the code to github111https://github.com/ictnlp/Multiscale-Contextualization.
@@ 1 Introduction
However, a feature of UTF-8 encoding hinders conventional Transformer model Vaswani et al. (2017) from adapting well to byte-based vocabulary: a single character may correspond to 1 to 4 UTF-8 bytes. The number is 1 for English characters, but Arabic and many Asian languages require multiple bytes to represent a single character. Therefore, sometimes a single byte does not have a determined meaning; it requires the integration of local information to encode its semantics. To address that, various methods have been proposed for integrating local contextual information. SU4MT Huang et al. (2023) learns contextual information with an Attentive Semantic Fusion layer, but requires accurate segmentation. MEGABYTE Yu et al. (2023) segments a sentence into blocks of 4 and simply concatenates the tokens. Charformer Tay et al. (2022) segments a sentence 4 times with block sizes ranging from 1 to 4 each, and employs mean-pooling to perform local integration. The weighted-sum of 4 results yields the final output. LOBEF Sreedhar et al. (2023) proposed Byte-nCF, replacing mean-pooling in Charformer with Convolutional Neural Networks (CNNs) for better performance.
@@ 3.1 Datasets
We use a multilingual TED corpus of 59 languages Qi et al. (2018), TED-59, which includes both high and low-resource languages. All cases are English-centered. We collect the raw data from Salesky et al. (2023) and preprocess it with two subword-level vocabularies and a byte-level one. For subword-based baseline system, we leverage SentencePiece Kudo and Richardson (2018) and train a 32k vocabulary on the training set. We also incorporate the 250k mBART-50 Liu et al. (2020); Tang et al. (2020) vocabulary for full lexical coverage. For Byte-level systems, we preprocess data with a 256 vocabulary using scripts222https://github.com/UriSha/EmbeddinglessNMT/blob/master/embeddingless_scripts/byte_preprocess.sh from Shaham and Levy (2021).
@@ 3.2 Models
Byte-nCF Sreedhar et al. (2023): A strong byte-based model performing well under low-resource settings. The structure hyper-parameters are of default setting333https://github.com/makeshn/LOBEF_Byte_NMT/blob/main/embeddingless_scripts/train_byte_ncf.sh.
@@ 4.4 Contextualization Scales
For the selection of series, which reflect the contextualization scales applied in a model, we experiment the small scales "0,0,1,1,3,3,5,5", the large scales "0,0,1,1,5,5,7,7", and the balanced scales "0,0,1,1,3,5,5,7". These models can leverage information of granularities of "1,3,5", "1,5,7", and "1,3,5,7" respectively444We have also experimented the granularities ”9”, ”11” and ”13”, but they are harmful to translation quality. It demonstrates the model is unable to process information from too many tokens well..
| small | large | balanced | |
| Byte-1 | 30.190* | 29.941 | 30.219 |
| Byte-2 | 21.251 | 21.601* | 21.545 |
| Byte-3 | 10.302 | 10.686* | 10.712 |
The results are exhibited in Table 4. First, the model of balanced scales performs the best averagely, because it is provided with contextualized information of more scales and has more options to choose. Second, if we ignore the balanced one and compare the other models, the performance is related to the language groups. The "∗" sign indicates the better performance between two models. For "Byte-1" group, a small scale is sufficient to model certain semantics, so the smaller scaled model performs better. For "Byte-3" groups, it requires larger contextual scales to form certain meanings, so the larger scaled model performs better.
These discoveries shade light on the selection of series, which is the series should be compatible with the language of input text.